OpenStudy (anonymous):

4. Which of the equations below is equivalent to; 18(19x+13y)=15 A. 2x+6y=15 B. 4x+6y=270 C. x+3y=15 D. 19x+13y=270

OpenStudy (anonymous):

18(19x + 13y) = 15 divide the whole equation by 3 18/3(19x + 13y) = 15/3 6(19x + 13y) = 5 6*19x + 6*13y = 5 114x + 78y = 5

OpenStudy (anonymous):

None of urs seem to match , is the question correct ?